# To Retrieve or Not to Retrieve? Uncertainty Detection for Dynamic Retrieval Augmented Generation

### v1

@@ Abstract.
retrieval augmented generation, retrieval, uncertainty detection, language models, interfaces
ccs: Computing methodologies Natural language processingccs: Information systems Information retrievalccs: Information systems Users and interactive retrieval
1. Introduction
Recently, Large Language Models (LLMs) like ChatGPT (OpenAI, 2023), Gemini (Team et al., 2023), and others are showing impressive strides in tasks across numerous benchmarks (Srivastava et al., 2023). This success has been largely owed to their exposure to massive training data and successive fine-tuning of instruction datasets. To increase the helpfulness and decrease the harmfulness of the models, they are being further fine-tuned over preference collections (Bai et al., 2022; Ouyang et al., 2022; Rafailov et al., 2024).
Further, Retrieval Augmented Generation (RAG) (Lewis et al., 2020; Dhole, 2024a; Dhole et al., 2024c), in the effort to mitigate hallucinations, enriches these models with domain-specific information and tackles scenarios where the intrinsic knowledge of the base model falls short. By integrating externally retrieved content during the generation phase, RAG enhances the model’s ability to produce less hallucinatory and domain-conditioned responses. This approach has been particularly valuable in complex applications such as long-form generation like multi-hop question answering, which often requires multiple retrievals to address a query comprehensively.
However, to optimize the efficiency of RAG, retrieval should only be invoked when necessary — also referred to as conditional retrieval. Previous conditional RAG setups have explored multiple paradigms like low token probabilities (Jiang et al., 2023), external classifiers (Wang et al., 2023), or low entity popularity (Mallen et al., 2023) as indicators of the LLMs’ knowledge gaps. However, most of these methods fall short in either approximating knowledge gaps of the LLMs or lacking the ability to invoke retrieval dynamically.
On the other hand, with the potential of LLMs to hallucinate, there has been an increasing interest in uncertainty detection methods to gauge LLMs’ confidence in their outputs (Fadeeva et al., 2023). Unlike traditional methods that rely on rigid heuristics or external classifiers, uncertainty detection leverages the inherent variability in LLM-generated responses to estimate confidence dynamically.
For instance, semantic sets-based UD approaches (Lin et al., [n. d.]) group responses based on meaning, and use the number of clusters to directly reflect the level of uncertainty — with greater variability signaling higher uncertainty. Similarly, spectral methods using eigenvalue Laplacians quantify response diversity by identifying strong or weak clustering patterns in pairwise similarity graphs. These approaches align with the probabilistic nature of LLMs as well as adaptively gauge uncertainty based on output coherence, making them more robust against adversarial or ambiguous inputs.
@@ 1. Introduction
In that regard, we resort to a conditional RAG system and employ numerous uncertainty detection metrics to test the need for invoking retrieval. Our RAG system performs forward-looking active retrieval in the style of Jiang et al. (2023) (Jiang et al., 2023).
@@
  • We design a retrieval augmented generation with dynamic retrieval
@@ 2. Related Work
3. Tasks and Datasets
We conduct experiments on the 2WikiMultihopQA dataset (Ho et al., 2020), a multi-hop open domain question answering (QA) dataset that tests the reasoning and inference skills of question-answering models. Questions in this dataset generally require two steps of reasoning to deduce the final answer, and the information for each step of reasoning can be obtained through referencing external information viz., Wikipedia passages.
4. Approach
@@
4.1. Uncertainty Evaluation of Future Sentence
Given a query $\mathbf{q}$, a retriever $\mathbf{R}$, a text generator $\mathbf{G}$, and a black box uncertainty estimation function $\mathbf{U}$, and partially generated sequence $t_{<i}$ until time step $i$, – we first generate a temporary sentence $t_{n}$ in the style of FLARE (Jiang et al., 2023).
@@ 4.1. Uncertainty Evaluation of Future Sentence
(1) $t_{i}=\mathbf{P}\{\mathbf{q},\ldots,y_{i-1}\}$
@@
(2) $y_{i}=\begin{cases}\mathbf{P}\{d_{1},\ldots,d_{k},\mathbf{q},\ldots,y_{i-1}\}&\text{if}~{}\mathbf{U}(t_{i})>\theta_{\mathbf{U}}\\ \mathbf{P}\{\mathbf{q},\ldots,y_{i-1}\}&otherwise\end{cases}$
where $d_{1}\ldots d_{k}$ are obtained from a retrieval system $\phi$.
(3) $d_{1}\ldots d_{k}:=\phi(\mathbf{q})$
4.2. Sequence Level Uncertainty Evaluation Measures
@@
    The degree matrix of the adjacency graph is also used to compute the uncertainty estimate (Lin et al., 2023). A node that is well-connected to other nodes, might be less uncertain. We use two similarity metrics for computing the degree matrix.
@@
Degree Matrix Jaccard U ¿ 0.4 Sub-Query 24 1.46 3.67 0.593
Eccentricity U ¿ 2 Sub-Query 22 2.23 4.05  0.605
Semantic Sets U ¿ 2 Sub-Query 23 2.52 4.09 0.411
Degree Matrix NLI U ¿ 0.5 Sub-Query 24 2.25 4.00 0.535
Table 1. Performance Metrics over a smaller seed set
@@
Degree Matrix Jaccard U ¿ 0.4 1.80 3.61 0.57 0.453 0.533 0.538
@@
Table 2. Performance Metrics for Different Uncertainty Estimators for 75 examples.
4.3. Subquery Generation for Retrieval
We resort to retrieving relevant knowledge to account for the information that the model is lacking to answer the question. FLARE (Jiang et al., 2023) generates a retrieval query for the missing entity in the temporary sentence by using the sentence with the low probability token removed or by prompting an external question generator to generate a question for the missing entity as the answer. We generalize this by instead prompting the model to generate a subquery to figure out the missing information needed to answer the user query in an open-ended manner.
We define a subquery generator $\mathbf{S_{Q}}$ which takes in as input few-shot exemplars of subqueries, the current user query $\mathbf{q}$, and the current partial answer sentences uttered in chain-of-thought (Wei et al., 2022) fashion. It seeks to generate subqueries to get a specific piece of information not generated in the partial answer sentences but is needed to answer $\mathbf{q}$. Once this subquery is generated, we use this subquery to retrieve additional passages from the external retriever $\mathbf{R}$. These passages are then appended to the user input, and the generation continues.
@@ 4.3. Subquery Generation for Retrieval
5. Setup
The generator used in all experiments was GPT-3 (davinci-002) (Brown et al., 2020), and the retriever employed was BM25 through PyTerrier (Macdonald et al., 2021; Dhole, 2024b). The base code used for conducting the experiments and computing the metrics presented in the tables was obtained from the active RAG setup by Jiang et al. (2023). For uncertainty detection, we resort to the Fadeeva et al. (2023)’s LM-Polygraph library.
@@ 6. Results
Triggering retrieval, when uncertainty computed through Eccentricity i.e. U ¿ 2, led to the highest F1 score of 0.605, with a lesser number of search operations. This approach balanced retrieval efficiency and task performance better than other methods. It required half the number of search operations than an Always Retrieve approach. Semantic Sets’ innovative clustering approach performed poorly, with an F1 score of 0.411. Using entailment-based similarity to compute uncertainty via the Degree Matrix NLI measure achieved an F1 score of 0.535, comparable to the baseline. The lightweight Degree Matrix (Jaccard) necessitated the least number of retrieval operations to perform better than an Always Retrieve baseline.
@@ 7. Conclusion
Besides, we feel that uncertainty detection might become more mainstream as the propensity for hallucination in LLMs increases, and as end applications demand more confidence and interpretability (Dhole et al., 2024c) in their outputs making uncertainty detection a necessity. Our work focuses on exploiting uncertainty detection for RAG, especially where retrieval can be expensive like the usage of heavy and composite retrieval systems employing numerous components like reformulation (Dhole et al., 2024b; Dhole and Agichtein, 2024; Dhole et al., 2024a), dense retrieval (Santhanam et al., 2021), reranking, etc.
8. Ethical Considerations
When evaluating large language models (LLMs), it is essential to adopt a sociotechnical perspective (Dhole, 2023), acknowledging that their outputs are influenced by both social contexts and technical design choices. Proper safeguards should be in place to mitigate biases and prevent the generation of harmful or toxic content. Furthermore, the uncertainty detection approaches we employed rely on estimations derived from various neural network computations, which are inherently shaped by the data on which the models are trained. Consequently, it is critical to thoroughly test uncertainty detection methods to ensure they meet the requirements of the intended applications.
@@ Acknowledgements
The author would like to thank Eugene Agichtein for his insightful discussions. This material is based upon work supported by a Microsoft Accelerating Foundation Models Research Award.

### v3

@@ Abstract
1 Introduction
Recently, Large Language Models (LLMs) like ChatGPT OpenAI (2023), Gemini Team et al. (2023), and others are showing impressive strides in tasks across numerous benchmarks Srivastava et al. (2023). This success has been largely owed to their exposure to massive training data and successive fine-tuning of instruction datasets. To increase the helpfulness and decrease the harmfulness of the models, they are being further fine-tuned over preference collections Bai et al. (2022); Ouyang et al. (2022); Rafailov et al. (2024).
Further, Retrieval Augmented Generation (RAG) Lewis et al. (2020); Dhole (2024a); Dhole et al. (2024), in the effort to mitigate hallucinations, enriches these models with domain-specific information and tackles scenarios where the intrinsic knowledge of the base model falls short. By integrating externally retrieved content during the generation phase, RAG enhances the model’s ability to produce less hallucinatory and domain-conditioned responses. This approach has been particularly valuable in complex applications such as long-form generation like multi-hop question answering, which often requires multiple retrievals to address a query comprehensively.
However, to optimize the efficiency of RAG, retrieval should only be invoked when necessary — also referred to as conditional retrieval. Previous conditional RAG setups have explored multiple paradigms like low token probabilities Jiang et al. (2023), external classifiers Wang et al. (2023), or low entity popularity Mallen et al. (2023) as indicators of the LLMs’ knowledge gaps. However, most of these methods fall short in either approximating knowledge gaps of the LLMs or lacking the ability to invoke retrieval dynamically.
On the other hand, with the potential of LLMs to hallucinate, there has been an increasing interest in uncertainty detection methods to gauge LLMs’ confidence in their outputs Fadeeva et al. (2023). Unlike traditional methods that rely on rigid heuristics or external classifiers, uncertainty detection leverages the inherent variability in LLM-generated responses to estimate confidence dynamically.
For instance, semantic sets-based UD approaches Lin et al. (2023) group responses based on meaning, and use the number of clusters to directly reflect the level of uncertainty — with greater variability signaling higher uncertainty. Similarly, spectral methods using eigenvalue Laplacians quantify response diversity by identifying strong or weak clustering patterns in pairwise similarity graphs. These approaches align with the probabilistic nature of LLMs as well as adaptively gauge uncertainty based on output coherence, making them more robust against adversarial or ambiguous inputs.
@@ 1 Introduction
In that regard, we resort to a conditional RAG system and employ numerous uncertainty detection metrics to test the need for invoking retrieval. Our RAG system performs forward-looking active retrieval in the style of Jiang et al. (2023).
@@
  • We design a method that performs retrieval augmented generation with dynamic retrieval through uncertainty detection
@@ 2 Related Work
Wang et al. (2024) proposed Self-DC that tackled compositional questions via iterative divide-and-conquer based on LLM certainty. Yao et al. (2024) propose utilising the model’s internal states to estimate uncertainty and deciding whether to retrieve or not.
@@
3 Tasks and Datasets
We conduct experiments on the 2WikiMultihopQA dataset Ho et al. (2020), a multi-hop open domain question answering (QA) dataset that tests the reasoning and inference skills of question-answering models. Questions in this dataset generally require two steps of reasoning to deduce the final answer, and the information for each step of reasoning can be obtained through referencing external information viz., Wikipedia passages.
4 Approach
@@
4.1 Uncertainty Evaluation of Future Sentence
Given a query $\mathbf{q}$, a retriever $\mathbf{R}$, a text generator $\mathbf{G}$, and a black box uncertainty estimation function $\mathbf{U}$, and partially generated sequence $t_{<i}$ until time step $i$, – we first generate a temporary sentence $t_{n}$ in the style of FLARE Jiang et al. (2023).
@@ 4.1 Uncertainty Evaluation of Future Sentence
$t_{i}=\mathbf{G}(\mathbf{P}\{\mathbf{q},\ldots,y_{i-1}\})$ (1)
@@
$y_{i}=\begin{cases}\mathbf{G}(\mathbf{P}\{d_{1},\ldots,d_{k},\mathbf{q},\ldots,y_{i-1}\})&\text{if}~{}\mathbf{U}(t_{i})>\theta_{\mathbf{U}}\\ \mathbf{G}(\mathbf{P}\{\mathbf{q},\ldots,y_{i-1}\})&otherwise\end{cases}$ (2)
where $d_{1}\ldots d_{k}$ are obtained from a retrieval system $\mathbf{R}$.
$d_{1}\ldots d_{k}:=\mathbf{R}(\mathbf{q})$ (3)
4.2 Sequence Level Uncertainty Evaluation Measures
@@
    The degree matrix of the adjacency graph is also used to compute the uncertainty estimate Lin et al. (2023). A node that is well-connected to other nodes, might be less uncertain. We use two similarity metrics for computing the degree matrix.
@@
Degree Matrix Jaccard U $>$ 0.4 Sub-Query 24 1.46 3.67 0.593
Eccentricity U $>$ 2 Sub-Query 22 2.23 4.05  0.605
Semantic Sets U $>$ 2 Sub-Query 23 2.52 4.09 0.411
Degree Matrix NLI U $>$ 0.5 Sub-Query 24 2.25 4.00 0.535
Table 1: Performance Metrics over a smaller seed set
@@
Degree Matrix Jaccard U $>$ 0.4 1.80 3.61 0.57 0.453 0.533 0.538
@@
Table 2: Performance Metrics for Different Uncertainty Estimators for 75 examples.
4.3 Subquery Generation for Retrieval
We resort to retrieving relevant knowledge to account for the information that the model is lacking to answer the question. FLARE Jiang et al. (2023) generates a retrieval query for the missing entity in the temporary sentence by using the sentence with the low probability token removed or by prompting an external question generator to generate a question for the missing entity as the answer. We generalize this by instead prompting the model to generate a subquery to figure out the missing information needed to answer the user query in an open-ended manner.
We define a subquery generator $\mathbf{S_{Q}}$ which takes in as input few-shot exemplars of subqueries, the current user query $\mathbf{q}$, and the current partial answer sentences uttered in chain-of-thought Wei et al. (2022) fashion. It seeks to generate subqueries to get a specific piece of information not generated in the partial answer sentences but is needed to answer $\mathbf{q}$. Once this subquery is generated, we use this subquery to retrieve additional passages from the external retriever $\mathbf{R}$. These passages are then appended to the user input, and the generation continues.
@@ 4.3 Subquery Generation for Retrieval
5 Setup
The generator used in all experiments was GPT-3 (davinci-002) Brown et al. (2020), and the retriever employed was BM25 through PyTerrier Macdonald et al. (2021); Dhole (2024b). The base code used for conducting the experiments and computing the metrics presented in the tables was obtained from the active RAG setup by Jiang et al. (2023). For uncertainty detection, we resort to the Fadeeva et al. (2023)’s LM-Polygraph library.
@@ 6 Results
Triggering retrieval, when uncertainty computed through Eccentricity i.e. $U>2$, led to the highest F1 score of 0.605, with a lesser number of search operations. This approach balanced retrieval efficiency and task performance better than other methods. It required half the number of search operations than an Always Retrieve approach. Semantic Sets’ innovative clustering approach performed poorly, with an F1 score of 0.411. Using entailment-based similarity to compute uncertainty via the Degree Matrix NLI measure achieved an F1 score of 0.535, comparable to the baseline. The lightweight Degree Matrix (Jaccard) necessitated the least number of retrieval operations to perform better than an Always Retrieve baseline.
@@ 7 Conclusion
Besides, we feel that uncertainty detection might become more mainstream as the propensity for hallucination in LLMs increases, and as end applications demand more confidence and interpretability Dhole et al. (2024) in their outputs making uncertainty detection a necessity. Our work focuses on exploiting uncertainty detection for RAG, especially where retrieval can be expensive like the usage of heavy and composite retrieval systems employing numerous components like reformulation, dense retrieval Santhanam et al. (2021), reranking, etc.
8 Ethical Considerations
When evaluating large language models (LLMs), it is essential to adopt a sociotechnical perspective Dhole (2023), acknowledging that their outputs are influenced by both social contexts and technical design choices. Proper safeguards should be in place to mitigate biases and prevent the generation of harmful or toxic content. Furthermore, the uncertainty detection approaches we employed rely on estimations derived from various neural network computations, which are inherently shaped by the data on which the models are trained. Consequently, it is critical to thoroughly test uncertainty detection methods to ensure they meet the requirements of the intended applications.
@@ Acknowledgements
The author would like to thank Eugene Agichtein for insightful discussions and the anonymous reviewers for their useful feedback. The author would also like to thank Microsoft for providing OpenAI credits through the Microsoft Accelerating Foundation Models Research Award.